# Fine-grained Attention I/O Complexity: Comprehensive Analysis for Backward Passes

### v1

@@
Fine-grained Attention I/O Complexity: Comprehensive Analysis for Backward Passes
Xiaoyu Li xli216@stevens.edu. Stevens Institute of Technology.    Yingyu Liang yingyul@hku.hk. The University of Hong Kong. yliang@cs.wisc.edu. University of Wisconsin-Madison.    Zhenmei Shi zhmeishi@cs.wisc.edu. University of Wisconsin-Madison.    Zhao Song magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the University of California, Berkeley.    Yufa Zhou yufazhou@seas.upenn.edu. University of Pennsylvania.
Large Language Models (LLMs) have demonstrated remarkable capabilities in processing long-context information. However, the quadratic complexity of attention computation with respect to sequence length poses significant computational challenges, and I/O aware algorithms have been proposed. This paper presents a comprehensive analysis of the I/O complexity for attention mechanisms, focusing on backward passes by categorizing into small and large cache scenarios. Using the red-blue pebble game framework, we establish tight bounds on I/O complexity across all cache sizes. We confirm that the de facto standard I/O aware algorithm FlashAttention is optimal for both forward and backward passes for the large cache size scenario. For small cache sizes, we provide an algorithm that improves over existing methods and achieves the tight bounds. Additionally, we extend our analysis to sparse attention, a mainstream speeding-up approach, deriving fine-grained lower bounds for both forward and backward passes and both small and large caches. Our findings complete the theoretical foundation for I/O complexity in attention mechanisms, offering insights for designing efficient algorithms of LLM training and inference.
@@ 1.1 Our Contributions
In this work, we analyze the I/O complexity in the same setting as the existing work of FlashAttention [26] and SY [102]. We consider a two-level memory hierarchy consisting of a small but fast layer called the cache and a large but slower layer referred to as memory. The I/O complexity quantifies the data transfer between these two layers, which can be formally defined as a red-blue pebble game [49] as in Definition 3.4. We study the exact attention computation using standard matrix multiplication as the existing work [^2] Note that there are many fast matrix multiplication methods. We do not study them, as they are hard to be parallelized. Standard matrix multiplication is still the most popular implementation on GPU, e.g., PyTorch. We refer readers to Section 3 for more details. and focus on backward gradient computation. We establish matching I/O complexity upper and lower bounds for attention backward computation (formalized in Theorem 1.1 and illustrated in Fig. 1). Combined with the attention forward results from SY [102], this completes the theory of I/O complexity in the attention mechanism.
Our main result is stated as follows:
@@
Let $n$ be the sequence length, $d$ the head dimension, and $M$ the cache size. The I/O complexity of attention backward computation under standard matrix multiplication is
$\Theta\left(\min\left\{\frac{n^{2}d^{2}+nd^{3}}{M},\frac{n^{2}d+nd^{2}}{\sqrt{M}}\right\}\right).$
@@
Therefore, we provide the optimal bounds and algorithms for backward passes for all cache sizes. This fully characterizes the I/O complexity of attention forward/backward when combined with existing results on forward passes [102]. Notably, we confirm that FlashAttention is optimal for both the forward and backward passes when the cache size is large enough $M=\Omega(d^{2})$.
Moreover, in recent years, sparse attention has become another mainstream method for speeding up the training process of transformer-based models [18, 122, 16]. These approaches mainly focus on techniques for sparsifying the attention matrix, thereby reducing the quadratic bottleneck in running time. However, it remains unknown whether this method can be integrated with I/O-aware algorithms like FlashAttention. Consequently, we further analyze the I/O complexity of sparse attention to provide theoretical guarantees, offering fine-grained lower bounds.
Let $Z_{\mathrm{input}}$ and $Z_{\mathrm{QK}}$ be the number of nonzero entries of the input matrix and the key-query matrix, respectively. Then any algorithm for both attention forward and backward computation using sparse semi-ring matrix multiplication has I/O complexity
Our I/O complexity lower bound for sparse attention recovers the lower bound for both attention forward and backward passes when matrices involved in attention computation are dense, i.e., $Z_{\mathrm{input}}=\Omega(nd),Z_{\mathrm{QK}}=\Omega(n^{2})$. In such case, our lower bound reads as $\Omega(\min\{\frac{n^{2}d^{2}}{M},\frac{n^{2}d}{\sqrt{M}}\})$, matching Theorem 1.1.
@@
Roadmap.
In Section 2, we review related literature. In Section 3, we introduce the definitions and background necessary for our study. We present our main results in Section 4 and discuss the techniques we employed in Section 5. Section 6 concludes our paper.
@@ 2 Related Work
Attention Computation Acceleration.
The quadratic time complexity of attention computation with respect to the length of the input sequence [108] poses significant computational challenges, especially for long sequences. Consequently, accelerating attention computation has become a crucial research area. From a theoretical standpoint, numerous works focus on approximating the attention matrix to accelerate computation [48, 8, 9, 68, 10, 71]. Experimental approaches involve modifying model architectures and optimizing implementations to accelerate inference. Methods such as Mamba [39, 27], Linearizing Transformers [120, 80], Hopfield Models [55, 110, 51, 116, 109, 46, 47] and PolySketchFormer [123, 61] aim to improve model performance and inference speed. System-level optimizations, such as FlashAttention [26, 25, 91] and block-wise parallel decoding [96], address bottlenecks in attention mechanisms and enhance inference speed through efficient implementation strategies. Collectively, these advancements contribute to making attention mechanisms more scalable and efficient, facilitating the deployment of large-scale language models. [95] accelerates inference by compressing the input text.
Learning with Bounded Memory and I/O Complexity.
A common memory model in computational systems is the two-level memory hierarchy. In this model, there are two layers of memory: a small but fast layer called the cache, and a large but slower layer called the memory. The I/O (input/output) complexity of an algorithm measures its efficiency based on the number of data transfer operations it performs between the cache and the memory. The early work of HK [49] formulated the I/O complexity mathematically using the language of graph theory. Learning with bounded memory has been studied in various fields in machine learning such as online learning [101, 84, 86], convex optimization [78, 20], active learning [50], attention computation [5], and continual learning [21, 36].
Sparse Attention.
Over the past few years, there has been extensive research on sparse Transformer/Attention models with weights pruning and inputs pruning, aimed at accelerating computation and training [119, 94, 16, 104, 43, 98, 92, 66, 34, 19]. In practice, the attention matrix is sparse, significantly reducing computational costs. Theoretical studies, such as [118], have demonstrated that sparse transformers are expressive enough and can achieve universal approximation properties.
@@ 3 Preliminary
In this work, we consider using a standard algorithm for matrix multiplication, which means that for any two matrices $A\in\mathbb{R}^{n_{1}\times d},B\in\mathbb{R}^{d\times n_{2}}$, each entry of $AB$ is computed by $(AB)_{i,j}=\sum_{k=1}^{d}A_{i,k}B_{k,j}$ for $i\in[n_{1}],j\in[n_{2}]$. Note that this setting is also used in FlashAttetnion [26] and SY [102]. Then, we introduce some key concepts needed for this paper.
@@ 3.1 Key Concept of Attention
Note that, we have $\mathsf{Softmax}(A_{1}XA_{2}^{\top})=D^{-1}\exp(A_{1}XA_{2}^{\top})\in\mathbb{R}^{n\times n}$, and usually we call it the attention matrix. The above definition is general and encompasses both self-attention and cross-attention mechanisms in Transformer architectures. Specifically, self-attention occurs when $A_{1}=A_{2}=A_{3}$, meaning that the queries, keys, and values are all derived from the same source. In contrast, cross-attention happens when $A_{2}=A_{3}$, indicating that the keys and values come from one source while the queries come from the other.
@@
Let $A_{1},A_{2}\in\mathbb{R}^{n\times d}$. Let $p(X)\in\mathbb{R}^{n\times n}$ be defined in Definition B.9 (see Fig. 2 for an illustration). Let $L(X)$ be some loss function. The attention backward gradient for $X\in\mathbb{R}^{d\times d}$ is:
$\frac{\mathrm{d}L(X)}{\mathrm{d}X}=A_{1}^{\top}p(X)A_{2}.$
@@
Figure 3: This diagram shows a summation tree with $d=2$ in the computational graph for the backward passes of attention using standard matrix multiplication. The orange and green nodes represent the input nodes of the level-$1$ summation tree. The brown nodes, along with the blue nodes (output from the level-$1$ summation tree), serve as inputs for the level-$2$ summation tree. The purple nodes represent the target output. When $d$ gets larger, the summation tree will expand with additional layers, where each new layer introduces intermediate nodes that represent the sums of pairs of nodes from the previous layer, i.e., there will be total $1+\log_{2}d$ layer in total.
@@ 4 Main Results
Here we discuss two implications of Theorem 1.1. First, through the fine-grained analysis, our result identifies a critical point at $M=d^{2}$, where the I/O complexity changes its behavior. For $M=o(d^{2})$, we establish better upper and lower bounds compared to existing results, demonstrating that FlashAttention is not optimal in this regime. Second, when $M=\Omega(d^{2})$, Theorem 1.1 provides a tighter lower bound than existing work using red-blue pebble game (Definition 3.4), offering insights of algorithm design.
Second, by combining the results of SY [102] with our findings, we provide a more general and tighter I/O complexity characterization of FlashAttention 1/2 [26, 25]. In the large cache scenario where $M=\Omega(d^{2})$, the attention forward I/O complexity is $\Theta(\frac{n^{2}d^{2}}{M})$, as discussed in Theorem 5.1 of SY [102]. Combining this result with our attention backward I/O complexity $\Theta(\frac{n^{2}d^{2}+nd^{3}}{M})$ (Theorem 1.1), we conclude that the overall complexity is $\Theta(\frac{n^{2}d^{2}+nd^{3}}{M})$. Thus, given the cache size is sufficiently large, i.e. $M=\Omega(d^{2})$, the I/O complexity of the forward and backward computation for FlashAttention 1/2 is optimal.
@@ 4.2 Small Cache
In the small cache case, we provide an upper bound below. Notice that this is better than the I/O complexity of FlashAttention which is $O(\frac{n^{2}d^{2}+nd^{3}}{M})>O(\frac{n^{2}d+nd^{2}}{\sqrt{M}})$ when $M=o(d^{2})$.
@@ 4.3 Lower Bound of Sparse Attention Forward and Backward Passes
Sparse attention is a generalization of standard attention and has been popular in practical applications. We refer readers to Section 2 for more discussion. To state our results, we first introduce some notations. For any matrix $A$, we use $\operatorname{nnz}(A)$ to denote the number of non-zero entries in the matrix $A$. We assume that sparse matrices are stored by listing only their non-zero entries along with their coordinates. We assume sparse semi-ring matrix multiplication, which restricts operations to addition and multiplication of these entries. Each output entry $(AB)_{i,j}$ can only be computed as the sum of products given by $\sum_{k}A_{i,k}B_{k,j}$.
@@
$\Omega\left(\min\left\{\frac{Z_{A}^{2}+Z_{A}Z_{X}}{M},\frac{Z_{A}\sqrt{Z_{AXA}}+\sqrt{Z_{A}Z_{X}Z_{AXA}}}{\sqrt{M}}\right\}\right).$
@@ 5 Technical Overview
Upper Bound of Large Cache. In Section D, we present algorithms for attention backward in the large cache case, where $M=\Omega(d^{2})$. Similar to FlashAttention, the $n\times n$ attention matrix $f$ (Definition B.5) cannot be directly loaded into cache, even though it has been computed and can be stored in memory. The overall algorithm (Algorithm 9) consists of two phases. In Phase 1 (Algorithm 7), we compute $S=A_{1}X$ and $h=A_{3}Y$, and these two matrices are then passed to Phase 2. In Phase 2 (Algorithm 8), the inputs are matrices $A_{1},A_{2},S,h,O,\mathrm{d}O\in\mathbb{R}^{n\times d}$ (Definitions 3.1, B.6, B.7, and B.8), and vector $l\in\mathbb{R}^{n}$ (Definition B.4). We vertically divide the inputs into row block matrices of size $B_{r}\times d$ or $B_{c}\times d$, where $B_{r}=\min\{\lceil M/4d\rceil,d\}$ and $B_{c}=\lceil M/4d\rceil$. Using these row block matrices as computation units, we follow the computational graph (Fig. 2) and FlashAttention’s procedure. After accounting for the reads and writes of the overall algorithm (Algorithm 9), we prove Theorem 4.1. Furthermore, when the cache size is as large as $\Theta(nd)$, the I/O complexity can be reduced to $O(nd+d^{2})$, which corresponds to the size of the input and output of the algorithm.
Lower Bound of Large Cache and Small Cache. In Section E, we establish the lower bounds for the I/O complexity of attention gradient computation in both large and small cache cases. Following Definitions 3.4 and 3.5, we analyze the red-blue pebble game on the computational graph of any attention backward algorithm using standard matrix multiplication. More specifically, the key concept is the $M$-partition, which decomposes the graph into subgraphs, ensuring that each subgraph satisfies conditions related to dominator and minimum sets (Definitions E.1, E.2, E.3, E.4, and E.5). Our proofs for the lower bound of backward passes builds upon the lemmas (Lemmas E.7 and E.8), which provide the foundation for relating the number of subgraphs to the I/O operations required. For the large cache scenario, $M=\Omega(d^{2})$, we demonstrate that the I/O complexity scales with the need to compute matrix products efficiently. In the small cache case, $M=o(d^{2})$, we show that higher I/O complexity is unavoidable due to the data transfers between cache and memory by reducing to the standard matrix multiplication. These analyses are formally established in the proofs of Theorems E.9 and E.10. In particular, our Theorems E.10, the small cache lower bound case, requires a new analysis deviation.
@@
The Softmax in Definition 3.1 can be changed to other non-linear activation functions and our lower bound still holds. It is because we must compute matrix multiplication of size $n\times d$ and $d\times n$ in non-linear attention. However, for linear attention, i.e., $A_{1}XA_{2}^{\top}A_{3}Y$, our lower bound is loose, since we can compute $\underbrace{A_{2}^{\top}}_{d\times n}\underbrace{A_{3}}_{n\times d}$ first, and then we have $\underbrace{A_{1}}_{n\times d}\underbrace{X}_{d\times d}\underbrace{A_{2}^{\top}A_{3}}_{d\times d}\underbrace{Y}_{d\times d}$.
Lower Bound of Sparse Attention Forward and Backward Passes. In Section F, we establish lower bounds on the I/O complexity of sparse attention computation for both forward and backward passes. Sparse matrix multiplication is considered, where only non-zero entries are stored and used in computations. We derive I/O complexity bounds based on the non-zero counts of input matrices and the I/O operations required for sparse matrix multiplication (Lemma F.1). We further extend these bounds to the matrix products involved in the attention mechanism (Lemma F.2), which requires multiple sparse matrix multiplication analysis. We analyze scenarios where matrices are stored in cache or require intermediate I/Os during computation to obtain the I/O complexity bounds for both forward and backward passes (Theorems F.3 and Theorem F.4), and Theorem 4.5 directly holds as a consequence.

### v2

# On Fine-Grained I/O Complexity of Attention Backward Passes

@@
On Fine-Grained I/O Complexity of Attention Backward Passes
Xiaoyu Li and Yingyu Liang and Zhenmei Shi and Zhao Song and Song Yue and Jiahao Zhang University of New South Wales. The University of Hong Kong. University of Wisconsin-Madison. magic.linuxkde@gmail.com. The Simons Institute for the Theory of Computing at the University of California, Berkeley. Northeastern University.
Large Language Models (LLMs) exhibit exceptional proficiency in handling extensive context windows in natural language. Nevertheless, the quadratic scaling of attention computation relative to sequence length creates substantial efficiency bottlenecks, necessitating the development of I/O-optimized algorithms. In this work, we conduct a systematic examination of the I/O complexity inherent in attention mechanisms, with a specific emphasis on the backward pass under both small and large cache settings. By leveraging the red-blue pebble game framework, we derive tight bounds for I/O complexity across the full spectrum of cache sizes. We validate that FlashAttention, one of the current industry standards, achieves optimality in the large-cache scenario for both forward and backward passes. Conversely, for small-cache environments, we introduce a novel algorithm that outperforms contemporary methods and successfully attains theoretical tight bounds. Furthermore, we expand our investigation to include sparse attention by establishing granular lower bounds for both forward and backward passes across all cache configurations. Ultimately, our results solidify the theoretical framework regarding I/O complexity in attention mechanisms, providing critical guidance for the development of efficient LLM training and inference systems.
@@ 1.1 Our Contributions
In this work, we analyze the I/O complexity in the same setting as the existing work of FlashAttention [27] and SY [104]. We consider a two-level memory hierarchy consisting of a small but fast layer called the cache and a large but slower layer referred to as memory. The I/O complexity quantifies the data transfer between these two layers, which can be formally defined as a red-blue pebble game [50] as in Definition 3.4. We study the exact attention computation using standard matrix multiplication as the existing work [^2] Note that there are many fast matrix multiplication methods. We do not study them, as they are hard to be parallelized. Standard matrix multiplication is still the most popular implementation on GPU, e.g., PyTorch. We refer readers to Section 3 for more details. and focus on backward gradient computation. We establish matching I/O complexity upper and lower bounds for attention backward computation (formalized in Theorem 1.1 and illustrated in Fig. 1). Combined with the attention forward results from SY [104], this completes the theory of I/O complexity in the attention mechanism. Our main result is stated as follows:
@@
Let $n$ be the sequence length, $d$ the head dimension, and $M$ the cache size. The I/O complexity of attention backward computation under standard matrix multiplication is $\Theta\left(\min\left\{\frac{n^{2}d^{2}+nd^{3}}{M},\frac{n^{2}d+nd^{2}}{\sqrt{M}}\right\}\right).$
@@
Therefore, we provide the optimal bounds and algorithms for backward passes for all cache sizes. This fully characterizes the I/O complexity of attention forward/backward when combined with existing results on forward passes [104]. Notably, we confirm that FlashAttention is optimal for both the forward and backward passes when the cache size is large enough $M=\Omega(d^{2})$. Moreover, in recent years, sparse attention has become another mainstream method for speeding up the training process of transformer-based models [18, 124, 16]. These approaches mainly focus on techniques for sparsifying the attention matrix, thereby reducing the quadratic bottleneck in running time. However, it remains unknown whether this method can be integrated with I/O-aware algorithms like FlashAttention. Consequently, we further analyze the I/O complexity of sparse attention to provide theoretical guarantees, offering fine-grained lower bounds.
Let $Z_{\mathrm{input}}$ and $Z_{\mathrm{QK}}$ be the number of nonzero entries of the input matrix and the key-query matrix, respectively. Then, any algorithm for both attention forward and backward computation using sparse semi-ring matrix multiplication has I/O complexity
Our I/O complexity lower bound for sparse attention recovers the lower bound for both attention forward and backward passes when matrices involved in attention computation are dense, i.e., $Z_{\mathrm{input}}=\Omega(nd),Z_{\mathrm{QK}}=\Omega(n^{2})$. In such case, our lower bound reads as $\Omega(\min\{\frac{n^{2}d^{2}}{M},\frac{n^{2}d}{\sqrt{M}}\})$, matching Theorem 1.1. The dividing point between small and large cache for sparse attention is $M=Z^{2}_{\mathrm{input}}/Z_{\mathrm{QK}}$, which also matches the dense case.
@@ 2 Related Work
Large Language Models. The exceptional success of generative large language models (LLMs), such as GPT-4 [2], Claude 3 [7], Gemini 1.5 [90], Llama 3.1 [74], Mistral Nemo [58], Phi 3.5 [4], is fundamentally attributed to the transformer architecture introduced by VSP+ [110] and all support at least 128k input token length. The transformer architecture and its self-attention mechanism have become indispensable in leading natural language processing (NLP) models [23], demonstrating remarkable capabilities across a diverse array of applications, including language translation [55], sentiment analysis [108], language modeling [77], the integration of differential privacy [91, 71], and multi-modal tasks [126, 72, 113]. Transformers’ emergent compositional abilities [31, 119] and proficiency in in-context learning [83, 75, 102] have led some to consider them as early indicators of Artificial General Intelligence (AGI) [14]. As such, the transformer architecture plays a pivotal role in advancing the field of AI.
Attention Computation Acceleration. The quadratic time complexity of attention computation with respect to the length of the input sequence [110] poses significant computational challenges, especially for long sequences. Consequently, accelerating attention computation has become a crucial research area. From a theoretical standpoint, numerous works focus on approximating the attention matrix to accelerate computation [49, 8, 9, 69, 10, 72]. Experimental approaches involve modifying model architectures and optimizing implementations to accelerate inference. Methods such as Mamba [40, 28], Linearizing Transformers [122, 81], Hopfield Models [56, 112, 52, 118, 111, 47, 48] and PolySketchFormer [125, 62] aim to improve model performance and inference speed. System-level optimizations, such as FlashAttention [27, 26, 92] and block-wise parallel decoding [97], address bottlenecks in attention mechanisms and enhance inference speed through efficient implementation strategies. Collectively, these advancements contribute to making attention mechanisms more scalable and efficient, facilitating the deployment of large-scale language models. [96] accelerates inference by compressing the input text.
Learning with Bounded Memory and I/O Complexity. A common memory model in computational systems is the two-level memory hierarchy. In this model, there are two layers of memory: a small but fast layer called the cache, and a large but slower layer called the memory. The I/O (input/output) complexity of an algorithm measures its efficiency based on the number of data transfer operations it performs between the cache and the memory. The early work of HK [50] formulated the I/O complexity mathematically using the language of graph theory. Learning with bounded memory has been studied in various fields in machine learning such as online learning [103, 85, 87], convex optimization [79, 20], active learning [51], attention computation [5], and continual learning [21, 37].
Due to space constraints, we move some related works to the Appendix A.
@@ 3 Preliminary
In this work, we consider using a standard algorithm for matrix multiplication, which means that for any two matrices $A\in\mathbb{R}^{n_{1}\times d},B\in\mathbb{R}^{d\times n_{2}}$, each entry of $AB$ is computed by $(AB)_{i,j}=\sum_{k=1}^{d}A_{i,k}B_{k,j}$ for $i\in[n_{1}],j\in[n_{2}]$. Note that this setting is also used in FlashAttetnion [27] and SY [104]. Although certain “fast” algorithms, such as Strassen [99] or Coppersmith–Winograd [22], have lower asymptotic computational complexity, they typically incur significantly large hidden constants and do not parallelize as efficiently on modern GPUs. In practice, libraries such as cuBLAS and CUTLASS optimized for standard GEMMs (General matrix multiplications), often outperform any known fast-matrix approach on the matrix sizes relevant to deep learning. Therefore, assuming standard matrix multiplication offers an accurate reflection of how attention computation is commonly carried out in real-world systems. Now, we introduce some key concepts needed for this paper.
@@ 3.1 Key Concept of Attention
Note that, we have
$\mathsf{Softmax}(A_{1}XA_{2}^{\top})=D^{-1}\exp(A_{1}XA_{2}^{\top})\in\mathbb{R}^{n\times n},$
and usually we call it the attention matrix. The above definition is general and encompasses both self-attention and cross-attention mechanisms in Transformer architectures. Specifically, self-attention occurs when $A_{1}=A_{2}=A_{3}$, meaning that the queries, keys, and values are all derived from the same source. In contrast, cross-attention happens when $A_{2}=A_{3}$, indicating that the keys and values come from one source while the queries come from the other.
@@
Let $A_{1},A_{2}\in\mathbb{R}^{n\times d}$. Let $p(X)\in\mathbb{R}^{n\times n}$ be defined in Definition B.9 (see Fig. 2 for an illustration). Let $L(X)$ be some loss function. The attention backward gradient for $X\in\mathbb{R}^{d\times d}$ is $\frac{\mathrm{d}L(X)}{\mathrm{d}X}=A_{1}^{\top}p(X)A_{2}.$
@@
Figure 3: This diagram shows a summation tree with $d=2$ in the computational graph for the backward passes of attention using standard matrix multiplication. The orange and green nodes represent the input nodes of the level-11 summation tree. The brown nodes, along with the blue nodes (output from the level-11 summation tree), serve as inputs for the level-22 summation tree. The purple nodes represent the target output. When $d$ gets larger, the summation tree will expand with additional layers, where each new layer introduces intermediate nodes that represent the sums of pairs of nodes from the previous layer, i.e., there will be a total $1+\log_{2}d$ layer in total.
@@ 4 Main Results
Here, we discuss two implications of Theorem 1.1. First, through the fine-grained analysis, our result identifies a critical point at $M=d^{2}$, where the I/O complexity changes its behavior. For $M=o(d^{2})$, we establish better upper and lower bounds compared to existing results, demonstrating that FlashAttention is not optimal in this regime. Second, when $M=\Omega(d^{2})$, Theorem 1.1 provides a tighter lower bound than existing work using red-blue pebble game (Definition 3.4), offering insights of algorithm design.
Moreover, by combining the results of SY [104] with our findings, we provide a more general and tighter I/O complexity characterization of FlashAttention 1/2 [27, 26]. In the large cache scenario where $M=\Omega(d^{2})$, the attention forward I/O complexity is $\Theta(\frac{n^{2}d^{2}}{M})$, as discussed in Theorem 5.1 of SY [104]. Combining this result with our attention backward I/O complexity $\Theta(\frac{n^{2}d^{2}+nd^{3}}{M})$ (Theorem 1.1), we conclude that the overall complexity is $\Theta(\frac{n^{2}d^{2}+nd^{3}}{M})$. Thus, given the cache size is sufficiently large, i.e., $M=\Omega(d^{2})$, the I/O complexity of the forward and backward computation for FlashAttention 1/2 is optimal.
@@ 4.2 Small Cache
In the small cache case, we provide an upper bound below. Notice that this is better than the I/O complexity of FlashAttention since $O(\frac{n^{2}d^{2}+nd^{3}}{M})>O(\frac{n^{2}d+nd^{2}}{\sqrt{M}})$ when $M=o(d^{2})$.
@@
Our theory suggests that, when the commonly used hidden dimension size increases while some commercial GPU cache sizes remain insufficiently large, our algorithm designed for small cache sizes would become relevant and useful. For example, the current network architectures usually set $d=128$, so the dividing point is approximately $d^{2}\times\text{size\_of(data type)}$, e.g., $16,384\times 32$-bit = $65.5$ KB for float32. For the NVIDIA A100 GPU, the size of each streaming multiprocessor (SM/L1 cache) is 192KB, so we can choose FlashAttention. However, for old GPUs such as NVIDIA GTX1060, the size of each SM is 48 KB, so the algorithm for the small cache size is suitable. Whether a cache is large or small depends on the relation between $M$ and $d$, not the absolute size. We can also view the small-cache case as a high-dimensional scenario, which may apply to other settings. Hence, our work provides theoretical insights and could guide future developments in attention mechanisms tailored to evolving hardware limitations.
@@ 4.3 Lower Bound of Sparse Attention Forward and Backward Passes
Sparse attention is a generalization of standard attention and has been popular in practical applications. We refer readers to Section 2 for more discussion. To state our results, we first introduce some notations. For any matrix $A$, we use $\operatorname{nnz}(A)$ to denote the number of non-zero entries in the matrix $A$. We assume that sparse matrices are stored by listing only their non-zero entries along with their coordinates. We assume sparse semi-ring matrix multiplication, which restricts operations to the addition and multiplication of these entries. Each output entry $(AB)_{i,j}$ can only be computed as the sum of products given by $\sum_{k}A_{i,k}B_{k,j}$.
@@
$\Omega\left(\min\left\{\frac{Z_{A}^{2}+Z_{A}Z_{X}}{M},\frac{Z_{A}\sqrt{Z_{AXA}}+\sqrt{Z_{A}Z_{X}Z_{AX}}}{\sqrt{M}}\right\}\right).$
@@ 5 Technical Overview
Upper Bound of Large Cache. In Section D, we present algorithms for attention backward in the large cache case, where $M=\Omega(d^{2})$. Similar to FlashAttention, the $n\times n$ attention matrix $f$ (Definition B.5) cannot be directly loaded into the cache, even though it has been computed and can be stored in memory. The overall algorithm (Algorithm 9) consists of two phases. In Phase 1 (Algorithm 7), we compute $S=A_{1}X$ and $h=A_{3}Y$, and these two matrices are then passed to Phase 2. In Phase 2 (Algorithm 8), the inputs are matrices $A_{1},A_{2},S,h,O,\mathrm{d}O\in\mathbb{R}^{n\times d}$ (Definitions 3.1, B.6, B.7, and B.8), and vector $l\in\mathbb{R}^{n}$ (Definition B.4). We vertically divide the inputs into row block matrices of size $B_{r}\times d$ or $B_{c}\times d$, where $B_{r}=\min\{\lceil M/4d\rceil,d\}$ and $B_{c}=\lceil M/4d\rceil$. Using these row block matrices as computation units, we follow the computational graph (Fig. 2) and FlashAttention’s procedure. After accounting for the reads and writes of the overall algorithm (Algorithm 9), we prove Theorem 4.1. When the cache size is as large as $\Theta(nd)$, the I/O complexity can be reduced to $O(nd+d^{2})$, which corresponds to the size of the input and output of the algorithm.
Lower Bound of Large Cache and Small Cache. In Section E, we establish the lower bounds for the I/O complexity of attention gradient computation in both large and small cache cases. Following Definitions 3.4 and 3.5, we analyze the red-blue pebble game on the computational graph of any attention backward algorithm using standard matrix multiplication. More specifically, the key concept is the $M$-partition, which decomposes the graph into subgraphs, ensuring that each subgraph satisfies conditions related to dominator and minimum sets (Definitions E.1, E.2, E.3, E.4, and E.5). Our proofs for the lower bound of backward passes builds upon the lemmas (Lemmas E.7 and E.8), which provide the foundation for relating the number of subgraphs to the I/O operations required. For the large cache scenario, $M=\Omega(d^{2})$, we demonstrate that the I/O complexity scales with the need to compute matrix products efficiently. In the small cache case, $M=o(d^{2})$, we show that higher I/O complexity is unavoidable due to the data transfers between cache and memory by reducing to the standard matrix multiplication. These analyses are formally established in the proofs of Theorems E.9 and E.10. Our Theorems E.10, the small cache lower bound case, requires a new analysis of deviation.
@@
The Softmax in Definition 3.1 can be changed to other non-linear activation functions and our lower bound still holds. It is because we must compute matrix multiplication of size $n\times d$ and $d\times n$ in non-linear attention. However, for linear attention, that is, $A_{1}XA_{2}^{\top}A_{3}Y$, our lower bound is loose. This is because we can compute $\underbrace{A_{2}^{\top}}_{d\times n}\underbrace{A_{3}}_{n\times d}$ first, and then compute $\underbrace{A_{1}}_{n\times d}\underbrace{X}_{d\times d}\underbrace{A_{2}^{\top}A_{3}}_{d\times d}\underbrace{Y}_{d\times d}$.
Lower Bound of Sparse Attention Forward and Backward Passes. In Section F, we establish lower bounds on the I/O complexity of sparse attention computation for both forward and backward passes. Sparse matrix multiplication is considered, where only non-zero entries are stored and used in computations. We derive I/O complexity bounds based on the non-zero counts of input matrices and the I/O operations required for sparse matrix multiplication (Lemma F.1). We extend these bounds to the matrix products involved in the attention mechanism (Lemma F.2), which requires multiple sparse matrix multiplication analysis. We analyze scenarios where matrices are stored in cache or require intermediate I/Os during computation to obtain the I/O complexity bounds for both forward and backward passes (Theorems F.3 and F.4), and Theorem 4.5 directly holds as a consequence.